Maxwell’s equations in integral form are composed of four equations

as follows:

1) Gauss’s law for electrical fields                                ;

2) Faraday’s law of electromagnetic induction                        ;

3) Gauss’s law for magnetic fields                                 ;

4) Ampere’s circuital law                                ;


1
Expert's answer
2021-06-28T06:08:02-0400

Part(1)

Gauss law for electric field


"\smallint E.dl=\frac{q}{\epsilon_0}"

Part(2)

Faraday laws for electromagnetic induction

"\smallint E. dl=-\frac{d}{dt}\smallint B.dl"

Part(3)

Gauss law for magnetic field


"\smallint B.da=0"

Part(4)

Ampear circuit law


"\smallint B.dl= \mu_0(I+\frac{d}{dt}(\epsilon_0\smallint E.dA))"